The temperature of 15.71 grams of gold rises from 32°C to 1,064°C, and then the gold melts completely. If gold’s specific heat is 0.1291 joules/gram degree Celsius and its heat of fusion is 63.5 joules/gram, how much energy is gained by the gold?

The gold gained a total of
joules of energy.
1
Expert's answer
2020-12-17T07:35:27-0500

"∆H=mass×specific heat capacity×∆T
=15.71g×(1064-32)°C×0.1291J\left(g^{\smash{-1}}\right)\left(°C^{\smash{-1}}\right)=2093.06J"


Enthalpy of fusion ="63.5\u00d715.71=997.585J"

Total energy ="2093.06+997.585=3090.64J"